# Computing Certificates of Strictly Positive Polynomials in Archimedean Quadratic Modules

### v1

@@ Abstract
New results on computing certificates of strictly positive polynomials in Archimedean quadratic modules are presented. The results build upon (i) Averkov’s method for generating a strictly positive polynomial for which a membership certificate can be more easily computed than the input polynomial whose certificate is being sought, and (ii) Lasserre’s method for generating a certificate by successively approximating a nonnegative polynomial by sums of squares. First, a fully constructive method based on Averkov’s result is given by providing details about the parameters; further, his result is extended to work on arbitrary subsets, in particular, the whole Euclidean space $\mathbb{R}^{n}$, producing globally strictly positive polynomials. Second, Lasserre’s method is integrated with the extended Averkov construction to generate certificates. Third, the methods have been implemented and their effectiveness is illustrated. Examples are given on which the existing software package RealCertify appears to struggle, whereas the proposed method succeeds in generating certificates. Several situations are identified where an Archimedean polynomial does not have to be explicitly included in a set of generators of an Archimedean quadratic module. Unlike other approaches for addressing the problem of computing certificates, the methods/approach presented is easier to understand as well as implement.
Key words: Quadratic module, certificate, sum of squares, Positivstellensatz
@@ 1 Introduction
Consider a polynomial set $G=\{g_{1},\ldots,g_{s}\}$ in the multivariate polynomial ring $\mathbb{R}[\bm{x}]$. The quadratic module generated by $G$ is $\operatorname{QM}(G)=\{\sigma_{0}+\sum_{i=1}^{s}\sigma_{i}\cdot g_{i}\,|\,\sigma_{i}\mbox{ is a sum of squares in }\mathbb{R}[\bm{x}]\mbox{ for }i\!=\!0,\ldots,s\}$ [10, 12]. Quadratic modules are fundamental objects in real algebraic geometry and are heavily related to Hilbert’s 17th problem [5] and Positivstellensatz. This paper addresses the problem of computing certificates that witness the membership of polynomials in quadratic modules in $\mathbb{R}[\bm{x}]$. The certificates here are also referred to as sum-of-squares representations in the literature. In particular, we present new results and constructive methods to compute the sum-of-squares multipliers, or certificates, of strictly positive polynomials in Archimedean finitely generated quadratic modules.
@@
Several contributions in the field identify conditions whenever representations in preorderings [^1] A preordering is a quadratic module closed under multiplication. and quadratic modules exist. Noteworthy are the seminal results by Schmüdgen [16] and Putinar’s Positivstellensatz [15] for strictly positive polynomials over compact semialgebraic sets. However, the above-mentioned results about membership problems do not produce certificates as their proofs are non-constructive. In the case of Hilbert’s 17th problem, as an example, for a given non-negative polynomial $p$ it is nontrivial to generate a certificate consisting of rational functions $\frac{a_{i}}{b_{i}}$’s such that $p=\sum_{i}\left({\frac{a_{i}}{b_{i}}}\right)^{2}$. A solution to the certificate problem for Archimedean preorderings is discussed in [17, 18] which is constructive by including an Archimedean polynomial in the set of generators; an algorithm is described using Gröbner basis techniques to lift positivity conditions to a simplex and using Pólya bounds [14] to compute certificates.
@@
Contributions. This paper presents new results about computing certificates of strictly positive polynomials in Archimedean quadratic modules. The results build upon (i) Averkov’s method [1] for generating a strictly positive polynomial for which a membership certificate can be more easily computed than the input polynomial whose certificate is being sought, and (ii) Lasserre’s method [6] for generating a certificate by successively approximating a non-negative polynomial by sums of squares.
@@
A fully constructive method based on Averkov’s result in [1] is given by providing details about the parameters left out there. Further, his result is extended to work on non-compact subsets, in particular, producing globally strictly positive polynomials on $\mathbb{R}^{n}$ (compare Theorem 3.3 and Lemma 2.3 for the details). This is discussed in Section 3. Then in Section 4 we extend Lasserre’s method to reduce any globally strictly positive polynomial by a single polynomial to a sum of squares. These two extensions of Averkov’s construction and Lasserre’s method are proved in a constructive manner and they serve as the theoretical and algorithmic foundations of our method for solving the certificate problem.
@@
To apply extended Averkov’s construction to the certificate problem, there are two technical problems to solve: construction of a polynomial in the quadratic module with a bounded non-negative set and a polynomial with a global lower bound. The former problem is easy when some generator of the quadratic module has a bounded non-negative set, otherwise we explicitly show how to construct such a polynomial in the univariate case and use an Archimedean polynomial in the multivariate case. Unlike the methods in the literature which depend on the inclusion of Archimedean polynomials in the generating sets, here we identify the situations where we can avoid inclusion of Archimedean polynomials and thus in these situations the certificates computed by our method are with respect to the original generators of the quadratic modules without the Archimedean polynomials.
@@
All the methods discussed in the paper have been implemented and their effectiveness is illustrated. Examples are given on which RealCertify, a software package for solving the certificate problem based on semidefinite programming [8], appears to struggle, whereas the proposed method succeeds in generating certificates. Several examples are given where an Archimedean polynomial does not have to be explicitly included in a given set of generators of an Archimedean quadratic module.
Related work. Among the literature most related to the topic of the paper, the following two publications stand out. In [9], the authors use the extended Gram matrix method to obtain an approximate certificate of the input polynomial in a quadratic module and a refinement step to compute sums of squares to represent the difference of the input polynomial and the approximate certificate (or the ”error term”). The method requires the polynomial set of generators to include an Archimedean polynomial as part of the assumptions. The authors also discuss time complexity for the termination of their algorithm and bounds of the degrees and bit size complexity of the coefficients in the certificates obtained; bounds of the degrees of the certificates rely on the complexity analysis discussed in [11]. Their implementation, RealCertify [8], relies on numerical methods, making them ineffective in certain situations. The performance of the approach proposed in the paper is compared with that of RealCertify later in Section 6.
More recently, [2] provides degree bounds on the sum-of-square multipliers in the certificate problem for a strictly positive polynomial $f$ in an Archimedean quadratic module $\operatorname{QM}(G)$. This is done in two steps. First, a new polynomial $h\in\operatorname{QM}(G)$ is constructed such that $f-h$ is strictly positive over the hypercube $\left[-1,1\right]^{n}$ by using a modified construction to Averkov’s result [1, Lemma 3.3]. Secondly, the construction in [7] is used to obtain a certificate of $f-h$ in the preordering generated by $1-\sum_{i=1}^{n}{x_{i}^{2}}$ that equals $\operatorname{QM}(1-\sum_{i=1}^{n}{x_{i}^{2}})$. This step relies on the Jackson kernel method, which seems highly complex and also difficult to algorithmize. Rearranging the previous results, the authors find a certificate of $f$ in $\operatorname{QM}(G\cup\{1-\sum_{i=1}^{n}{x_{i}^{2}}\})$. This approach also relies on adding an Archimedean polynomial to the generators. The authors do not report any implementation of their approach.
@@ 2 Preliminaries
2.1 Quadratic modules in $\mathbb{R}[x_{1},\ldots,x_{n}]$
Consider the multivariate polynomial ring $\mathbb{R}[x_{1},\ldots,x_{n}]$ over the field $\mathbb{R}$ of real numbers. We denote $\bm{x}=(x_{1},\ldots,x_{n})$, $\mathbb{R}[\bm{x}]=\mathbb{R}[x_{1},\ldots,x_{n}]$, and $\|\bm{x}\|=\sqrt{x_{1}^{2}+\cdots+x_{n}^{2}}$. For any $\bm{\alpha}=(a_{1},\ldots,a_{n})\in\mathbb{N}^{n}$, denote $\bm{x}^{\bm{\alpha}}=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$.
A polynomial in $\mathbb{R}[\bm{x}]$ is called a sum of squares if it can be expressed as a sum of squares of polynomials in $\mathbb{R}[\bm{x}]$. We denote the set of all sums of squares in $\mathbb{R}[\bm{x}]$ by $\sum\mathbb{R}[\bm{x}]^{2}$. Any sum of squares in $\mathbb{R}[\bm{x}]$ is a polynomial of even total degree and is non-negative over $\mathbb{R}^{n}$.
@@ 2.1 Quadratic modules in $\mathbb{R}[x_{1},\ldots,x_{n}]$
A polynomial set $M\subseteq\mathbb{R}[\bm{x}]$ is called a quadratic module in $\mathbb{R}[\bm{x}]$ if $M$ is closed under addition, $1\in M$, and for any $f\in\mathbb{R}[\bm{x}]$ and $m\in M$, $f^{2}m\in\mathbb{R}[\bm{x}]$. A quadratic module $M$ in $\mathbb{R}[\bm{x}]$ is finitely generated if there exists a finite set $G=\{g_{1},\dots,g_{s}\}\subseteq\mathbb{R}[\bm{x}]$ such that $M=\{\sigma_{0}+\sum_{i=1}^{s}\sigma_{i}g_{i}\,|\,\sigma_{i}\in\sum\mathbb{R}[\bm{x}]^{2},i\!=\!0,\ldots,s\}$. In this case, $M$ is said to be generated by $G$ and written as $M=\operatorname{QM}(G)$.
Let $G$ be a polynomial set in $\mathbb{R}[\bm{x}]$. Then the non-negative set of $G$, denoted by $\operatorname{S}(G)$, is defined to be $\{\bar{\bm{x}}\in\mathbb{R}^{n}\,|\,g_{i}(\bar{\bm{x}})\geq 0,i=1,\ldots,s\}$. A finitely generated quadratic module $\operatorname{QM}(G)$ in $\mathbb{R}[\bm{x}]$ is said to be bounded if the non-negative set $\operatorname{S}(G)$ is a bounded set in $\mathbb{R}^{n}$ and unbounded otherwise. Furthermore, a quadratic module in $\mathbb{R}[\bm{x}]$ is said to be Archimedean if it contains the polynomial $N-\sum_{i=1}^{n}x_{i}^{2}$ for some $N\in\mathbb{N}$, which is called Archimedean polynomial. It is clear that Archimedean quadratic modules are bounded.
Given a quadratic module $\operatorname{QM}(G)$ in $\mathbb{R}[\bm{x}]$, the fundamental membership problem is to determine whether an arbitrary polynomial $f$ belongs to $\operatorname{QM}(G)$, and this problem has been solved in specific cases. The well-known Putinar’s Positivstellensatz about strictly positive polynomials is recalled below.
Theorem 2.2 (Putinar Positivstellensatz, [15]).
Let $,G=\{g_{1},\ldots$$,g_{s}\}$ be a polynomial set in $\mathbb{R}[\bm{x}]$ such that $\operatorname{QM}(G)$ is Archimedean and $f\in\mathbb{R}[\bm{x}]$ be a polynomial that is strictly positive on $\operatorname{S}(G)$. Then $f\in\operatorname{QM}(G)$.
@@ 2.2 Averkov’s Lemma to aid computing certificates of Positivstellensatz theorems
Several (semi)constructive approaches exist in the literature to compute certificates of strictly positive polynomials in quadratic modules [19, 11, 1]. A common step in these approaches is the perturbation of the input polynomial $f\in\operatorname{QM}(G)$ by translating the positivity of $f$ over $\operatorname{S}(G)$ to the positivity of $f-h$ in some compact set $B$ where $h\in\operatorname{QM}(G)$. The following result due to Averkov is used in our construction. For self-containedness, the proof in [1] is sketched without giving details of the derivation.
Lemma 2.3 ([1, Lemma 3.3]).
Let $G=\{g_{1},\ldots,g_{s}\}$ be a polynomial set in $\mathbb{R}[\bm{x}]$ and $f$ be a polynomial in $\mathbb{R}[\bm{x}]$ such that $f>0$ on $\operatorname{S}(G)$. Then, for any compact subset $B$ of $\mathbb{R}^{n}$, a polynomial $h\in\operatorname{QM}(G)$ exists such that $f-h>0$ on $B$.
@@
Let $T=\{\bar{\bm{x}}\in B\leavevmode\nobreak\ |\leavevmode\nobreak\ f(\bar{\bm{x}})<0\}$. We use $a$ as the function $\bm{x}\mapsto(g_{1}(\bm{x}),\ldots,g_{s}(\bm{x}))$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{s}$. Since $a(B)$ is compact, there exists $\gamma>0$ such that $a(B)\subseteq(-\infty,2\gamma]^{s}$. By the assumption on $f$, $a(T)\cap[0,2\gamma]^{s}=\emptyset$. As $a(T)$ and $[0,2\gamma]^{s}$ are compact, there exists $\epsilon>0$ such that $a(T)\cap[-2\epsilon,\gamma]^{s}=\emptyset$. Thus, if $\bar{\bm{x}}\in B$ and $g_{i}(\bar{\bm{x}})\geq-2\epsilon$ for each $i=1,\dots,s$, then $f(\bar{\bm{x}})>0$. Consequently, $\mu:=\min\{f(\bar{\bm{x}})\leavevmode\nobreak\ |\leavevmode\nobreak\ \bar{\bm{x}}\in B\ \mbox{and}\ g_{i}(\bar{\bm{x}})\geq-2\epsilon\mbox{ for }i=1,\ldots,s\}>0$.
@@
$h=\sum_{i=1}^{s}g_{i}\left(\frac{g_{i}-\gamma}{\gamma+\epsilon}\right)^{2N}$ (1)
and denote
@@
For $\bar{\bm{x}}\in B$, if $g_{i}(\bar{\bm{x}})\geq-2\epsilon$ for each $i=1\ldots,s$, we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq\mu-sc(N)$; otherwise if $g_{i}(\bar{\bm{x}})<-2\epsilon$ for some $i\leavevmode\nobreak\ (1\leq i\leq s)$, we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq\min_{y\in B}f(y)-sc(N)+C(N)$. Since $c(N)\rightarrow 0$ and $C(N)\rightarrow+\infty$ as $N\rightarrow+\infty$, by choosing $N$ sufficiently large, we deduce $f(\bar{\bm{x}})-h(\bar{\bm{x}})>0$ for every $\bar{\bm{x}}\in B$. ∎

### v2

@@ Abstract
This paper addresses the problem of computing certificates that witness the membership of strictly positive polynomials in multivariate Archimedean quadratic modules. New algorithms for computing such certificates are presented based on (i) Averkov’s method for generating a strictly positive polynomial for which a certificate can be more easily computed, and (ii) Lasserre’s method for generating a certificate by successively approximating a non-negative polynomial by sums of squares.
A constructive method based on Averkov’s result is given by providing details about the parameters; further, his result is extended to work on arbitrary subsets, in particular the whole Euclidean space $\mathbb{R}^{n}$, producing globally strictly positive polynomials. Then Lasserre’s method is extended for monogenic Archimedean quadratic modules and integrated with the extended Averkov construction above to generate certificates.
All the presented methods have been implemented and their effectiveness is illustrated. Examples are given on which the existing software package RealCertify appears to struggle, whereas the proposed method succeeds in generating certificates. Several situations are identified where an Archimedean polynomial does not have to be explicitly included in a set of generators of an Archimedean quadratic module. Unlike other approaches for addressing the problem of computing certificates, the methods presented are easier to understand as well as to implement.
keywords:
Quadratic module , certificate , sum of squares , Positivstellensatz
[…]
@@ 1 Introduction
Consider a finite polynomial set $G=\{g_{1},\ldots,g_{s}\}$ in the multivariate polynomial ring $\mathbb{R}[X]$. The quadratic module generated by $G$ is $\operatorname{QM}(G)=\{\sigma_{0}+\sum_{i=1}^{s}\sigma_{i}\cdot g_{i}\,|\,\sigma_{i}$ is a sum of squares in $\mathbb{R}[X]\mbox{ for }i\!=\!0,\ldots,s\}$ Marshall (2008); Powers (2011). Quadratic modules are fundamental objects in real algebraic geometry and are heavily related to Hilbert’s (1888) 17th problem and Positivstellensatz. This paper addresses the problem of computing certificates that witness the membership of polynomials in quadratic modules in $\mathbb{R}[X]$. The certificates here are also referred to as sum-of-squares representations in the literature. In particular, we present new results and constructive methods to compute the sum-of-squares multipliers, or certificates, of strictly positive polynomials in Archimedean finitely generated quadratic modules.
@@
Several contributions in the field identify conditions whenever representations in preorderings [^1] A preordering is a quadratic module closed under multiplication. and quadratic modules exist. Noteworthy are the seminal results by Schmüdgen’s (1991) and Putinar’s (1993) Positivstellensatz for strictly positive polynomials over compact semialgebraic sets. However, the above-mentioned results about membership problems do not produce certificates as their proofs are non-constructive. In the case of Hilbert’s 17th problem, as an example, for a given non-negative polynomial $p$ it is nontrivial to generate a certificate consisting of rational functions $\frac{a_{i}}{b_{i}}$’s such that $p=\sum_{i}\left({\frac{a_{i}}{b_{i}}}\right)^{2}$. A constructive solution to the certificate problem for Archimedean preorderings is discussed in Schweighofer (2002, 2004) by including an Archimedean polynomial  [^2] Given an Archimedean quadratic module $Q$, an Archimedean polynomial of $Q$ is a polynomial of the form $N-\sum_{i=1}^{n}{x_{i}^{2}}$ with $N\in\mathbb{N}$ that belongs to $Q$. in the set of generators; an algorithm is described using Gröbner basis techniques to lift positivity conditions to a simplex and using Pólya bounds to compute certificates in Powers and Reznick (2001).
@@
Contributions. This paper presents new results about computing certificates of strictly positive polynomials in Archimedean quadratic modules. The results build upon (i) Averkov’s (2013) method for generating a strictly positive polynomial over a given compact subset $B$ of $\mathbb{R}^{n}$ for which a membership certificate can be more easily computed than the input polynomial whose certificate is being sought, and (ii) Lasserre’s (2006) method for generating a certificate by successively approximating a non-negative polynomial by sums of squares.
@@
A constructive method based on Averkov’s 2013 result is given by providing details about the parameters left out there. Further, his result is extended to work on any subsets, in particular, on $\mathbb{R}^{n}$ (compare ?THM? LABEL:lemma:d and ?THM? LABEL:lemma:c for details). These are discussed in Section 3. Then in Section 4 we extend Lasserre’s method to reduce any globally strictly positive polynomial by a single polynomial to a sum of squares. These two extensions of Averkov’s construction and Lasserre’s method are proved in a constructive manner, and they serve as the theoretical and algorithmic foundations of our method for solving the certificate problem. As a by-product, using these two extensions we give a new constructive proof of Putinar’s criterion for a quadratic module to be Archimedean (Theorem LABEL:thm:archimedean) in Section 5.
@@
To apply the extended Averkov’s construction to the certificate problem, there are two technical problems to solve: construction of a polynomial in the quadratic module with a bounded non-negative set and a polynomial with a global lower bound. The former problem is easy when some generator of the quadratic module has a bounded non-negative set, otherwise we explicitly show how to construct such a polynomial in the univariate case and use an Archimedean polynomial in the multivariate case. Unlike the methods in the literature which depend on the inclusion of Archimedean polynomials in the generating sets, we identify the situations where we can avoid inclusion of Archimedean polynomials and thus in these situations the certificates computed by our method are with respect to the original generators of the quadratic modules without the Archimedean polynomials.
@@
All the methods discussed in the paper have been implemented and their effectiveness is illustrated. Examples are given on which RealCertify, a software package for solving the certificate problem based on semidefinite programming in Magron and Safey El Din (2018b), appears to struggle, whereas the proposed method succeeds in generating certificates. Several examples are given where an Archimedean polynomial does not have to be explicitly included in an input set of generators of an Archimedean quadratic module.
Related work. Among the literature most related to the topic of the paper, the following two publications stand out. In Magron and Safey El Din (2018a), the authors use the extended Gram matrix method to obtain an approximate certificate of the input polynomial in a quadratic module and a refinement step to compute sums of squares to represent the difference of the input polynomial and the approximate certificate (or the “error term”). The method requires the polynomial set of generators to include an Archimedean polynomial as part of the assumptions. The authors also discuss time complexity for the termination of their algorithm and bounds of the degrees and bit size complexity of the coefficients in the certificates obtained; bounds of the degrees of the certificates rely on the complexity analysis discussed in Nie and Schweighofer (2007). Their implementation, RealCertify Magron and Safey El Din (2018b), relies on numerical methods, making them ineffective in certain situations. The performance of the approach proposed in the paper is compared with that of RealCertify later in Section 7.
More recently, Baldi and Mourrain (2023) provides degree bounds on the sum-of-square multipliers in the certificate problem for a strictly positive polynomial $f$ in an Archimedean quadratic module $\operatorname{QM}(G)$. This is done in two steps. First, a new polynomial $h\in\operatorname{QM}(G)$ is constructed such that fhf-h is strictly positive over the hypercube $\left[-1,1\right]^{n}$ by using a modified construction to Averkov’s (2013, [Lemma 3.3]) result. Secondly, the construction in Laurent and Slot (2022) is used to obtain a certificate of fhf-h in the preordering generated by $1-\sum_{i=1}^{n}{x_{i}^{2}}$ that equals $\operatorname{QM}(1-\sum_{i=1}^{n}{x_{i}^{2}})$. This step relies on the Jackson kernel method, which seems highly complex and also difficult to algorithmize. Rearranging the previous results, the authors find a certificate of $f$ in $\operatorname{QM}(G\cup\{1-\sum_{i=1}^{n}{x_{i}^{2}}\})$. This approach also relies on adding an Archimedean polynomial to the generators. The authors do not report any implementation of their approach.
@@ 2 Preliminaries
Throughout this paper, we use $\mathbb{R}$ to denote the field of real numbers, and $\mathbb{K}$ is a computable subfield of $\mathbb{R}$. Here, a computable field is a field for which there is an algorithm for computation of signs of elements of $\mathbb{K}$ Lachlan and Madison (1970); Strzebonski (2008). When talking about effective algorithms in this paper, we restrict ourselves to polynomials over $\mathbb{K}$ so that all the construction in the algorithms is feasible. For example, the field $\mathbb{Q}$ of rational numbers, a computable subfield of $\mathbb{R}$, is frequently used as the base field in the case of computation in real algebraic geometry.
2.1 Quadratic modules
Consider the multivariate polynomial ring $\mathbb{R}[X]=\mathbb{R}[x_{1},\ldots,x_{n}]$. We denote $X=(x_{1},\ldots,x_{n})$ and $\|X\|=\sqrt{x_{1}^{2}+\cdots+x_{n}^{2}}$. For any $\bm{\alpha}=(a_{1},\ldots,$ $a_{n})\in\mathbb{N}^{n}$, where $\mathbb{N}$ is the set of natural numbers, denote $X^{\bm{\alpha}}=x_{1}^{a_{1}}\cdots x_{n}^{a_{n}}$. We also use $\bar{\bm{x}}$ to denote a point in $\mathbb{R}^{n}$.
A polynomial in $\mathbb{R}[X]$ is called a sum of squares if it can be expressed as a sum of squares of polynomials in $\mathbb{R}[X]$. We denote the set of all sums of squares in $\mathbb{R}[X]$ by $\sum\mathbb{R}[X]^{2}$. Any sum of squares in $\mathbb{R}[X]$ is a polynomial of even total degree and is non-negative over $\mathbb{R}^{n}$.
@@ 2.1 Quadratic modules
A polynomial set $M\subseteq\mathbb{R}[X]$ is called a quadratic module in $\mathbb{R}[X]$ if $M$ is closed under addition, $1\in M$, and for any $f\in\mathbb{R}[X]$ and $m\in M$, $f^{2}m\in M$. A quadratic module $M$ in $\mathbb{R}[X]$ is finitely generated if there exists a finite set $G=\{g_{1},\dots,g_{s}\}\subseteq\mathbb{R}[X]$ such that $M=\{\sigma_{0}+\sum_{i=1}^{s}\sigma_{i}\cdot g_{i}\,|\,\sigma_{i}\in\sum\mathbb{R}[X]^{2},i\!=\!0,\ldots,s\}$. In this case, $M$ is said to be generated by $G$ and written as $M=\operatorname{QM}(G)$.
[…]
Theorem 2.2 (Putinar’s criterion, Putinar (1993)).
For a quadratic module $\operatorname{QM}(G)$ of $\mathbb{R}[X]$ with $G=\{g_{1},\ldots,g_{s}\}$, $\operatorname{QM}(G)$ is Archimedean if and only if there exists some $g\in\operatorname{QM}(G)$ such that $\operatorname{S}(g)$ is compact.
Given a quadratic module $\operatorname{QM}(G)$ in $\mathbb{R}[X]$, the fundamental membership problem is to determine whether an arbitrary polynomial $f$ belongs to $\operatorname{QM}(G)$, and this problem has been solved in specific cases. The well-known Putinar’s Positivstellensatz about strictly positive polynomials is recalled below.
Theorem 2.3 (Putinar Positivstellensatz, Putinar (1993)).
Let $G=\{g_{1},\ldots$ $g_{s}\}$ be a polynomial set in $\mathbb{R}[X]$ such that $\operatorname{QM}(G)$ is Archimedean and $f\in\mathbb{R}[X]$ be a polynomial that is strictly positive on $\operatorname{S}(G)$. Then $f\in\operatorname{QM}(G)$.
[…]
@@ 2.2 Averkov’s Lemma to aid computing certificates of Positivstellensatz theorems
Several (semi-)constructive approaches exist in the literature to compute or construct certificates of strictly positive polynomials in quadratic modules Schweighofer (2005); Nie and Schweighofer (2007); Averkov (2013). A common step in these approaches is the perturbation of the input polynomial $f\in\operatorname{QM}(G)$ by translating the positivity of $f$ over $\operatorname{S}(G)$ to the positivity of fhf-h in some compact set $B$ where $h\in\operatorname{QM}(G)$. In the following theorems, we restrict to the case $B\neq\emptyset$. When $B=\emptyset$, the statement holds trivially and no further construction is needed. The following result due to Averkov is used in our construction. For self-containedness, the proof in Averkov (2013) is sketched without giving details of the derivation.
Lemma 2.4 (Averkov (2013, Lemma 3.3)).
Let $G=\{g_{1},\ldots,g_{s}\}$ be a polynomial set in $\mathbb{R}[X]$ and $f$ be a polynomial in $\mathbb{R}[X]$ such that $f>0$ on $\operatorname{S}(G)$. Then, for any compact subset $B$ of $\mathbb{R}^{n}$, a polynomial $h\in\operatorname{QM}(G)$ exists such that fh>0f-h>0 on $B$.
@@
Let $T=\{\bar{\bm{x}}\in B\penalty 10000\ |\penalty 10000\ f(\bar{\bm{x}})\leq 0\}$. We use $a$ as the function $\bar{\bm{x}}\mapsto(g_{1}(\bar{\bm{x}}),\ldots,g_{s}(\bar{\bm{x}}))$ from $\mathbb{R}^{n}$ to $\mathbb{R}^{s}$. Since $a(B)$ is compact, there exists $\gamma>0$ such that $a(B)\subseteq(-\infty,2\gamma]^{s}$. By the assumption on $f$, $a(T)\cap[0,2\gamma]^{s}=\emptyset$. As $a(T)$ and $0,2\gamma]^{s}$ are compact, there exists $\epsilon>0$ such that $a(T)\cap[-2\epsilon,\gamma]^{s}=\emptyset$. Thus, if $\bar{\bm{x}}\in B$ and $g_{i}(\bar{\bm{x}})\geq-2\epsilon$ for each $i=1,\dots,s$, then $f(\bar{\bm{x}})>0$. Consequently, $\mu:=\min\{f(\bar{\bm{x}})\penalty 10000\ |\penalty 10000\ \bar{\bm{x}}\in B\ \mbox{and}\ g_{i}(\bar{\bm{x}})\geq-2\epsilon\mbox{ for }i=1,\ldots,s\}>0$.
@@
$h:=\sum_{i=1}^{s}\left(\frac{g_{i}-\gamma}{\gamma+\epsilon}\right)^{2N}\cdot g_{i},$ (1)
where $N\in\mathbb{N}$ and denote
@@
Then for any $\bar{\bm{x}}\in B$, if $g_{i}(\bar{\bm{x}})\geq-2\epsilon$ for each $i=1\ldots,s$, we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq\mu-sc(N)$; otherwise if $g_{i}(\bar{\bm{x}})<-2\epsilon$ for some $i\penalty 10000\ (1\leq i\leq s)$, we have $f(\bar{\bm{x}})-h(\bar{\bm{x}})\geq\min_{\bar{\bm{x}}\in B}f(\bar{\bm{x}})-sc(N)+C(N)$. Since $c(N)\rightarrow 0$ and $C(N)\rightarrow+\infty$ as $N\rightarrow+\infty$, by choosing $N$ sufficiently large, we deduce $f(\bar{\bm{x}})-h(\bar{\bm{x}})>0$ for every $\bar{\bm{x}}\in B$. ∎